View previous attempt Item 3 Item 3 Time Remaining 1 hour 59 minutes 5 seconds 01:59:05 Comans Corporation has two production departments, Milling and Customizing. The company uses a job-order costing system and computes a predetermined overhead rate in each production department. The Milling Department’s predetermined overhead rate is based on machine-hours and the Customizing Department’s predetermined overhead rate is based on direct labor-hours. At the beginning of the current year, the company had made the following estimates: Milling Customizing Machine-hours 15,000 20,000 Direct labor-hours 11,000 5,000 Total fixed manufacturing overhead cost $82,500 $29,500 Variable manufacturing overhead per machine-hour $ 2.00 Variable manufacturing overhead per direct labor-hour $ 4.10 During the current month the company started and finished Job A319. The following data were recorded for this job: Job A319: Milling Customizing Machine-hours 60 10 Direct labor-hours 50 40 Direct materials $690 $220 Direct labor cost $740 $660 If the company marks up its manufacturing costs by 10% then the selling price for Job A319 would be closest to: (Round your intermediate calculations to 2 decimal places.)
Answer #1

Milling overhead rate = (82500/15000+2) = 7.50 per MH

Customizing overhead rate = (29500/5000+4.1) = 10 per DLH

Calculate Selling price

Job A319
Direct material (690+220) 910
Direct labor (740+660) 1400
Milling overhead (60*7.5) 450
Customizing overhead (40*10) 400
Total manufacturing cost 3160
Markup 316
Selling price 3476